OpenStudy (y1066380):

How to solve d/dx((x+2)sqrt(x-1))?

OpenStudy (gorv):

\[\frac{ d ((x+2) \sqrt{x-1} )}{ dx }\]

OpenStudy (gorv):

\[\sqrt{x-1 } * \frac{ d(x+2) }{ dx }+(x+2)*\frac{ d \sqrt{x-1} }{ dx }\]

OpenStudy (gorv):

@y1066380 ...solve further now its normal diff.....

OpenStudy (y1066380):

The answer is

OpenStudy (y1066380):

I think @csmath is correct

OpenStudy (y1066380):

Thank you!!

OpenStudy (y1066380):

(Y) I know now. Thanks